OpenStudy (anonymous):

PROBABILITY QUESTION, HELP! The following table shows the probability distribution for a discrete random variable. X 23 25 26 31 34 38 P(X) 0.16 0.09 0.18 0.12 0.24 0.21 What is the mean of this discrete random variable? That is, what is E(X), the expected value of X?

hartnn (hartnn):

Expected value of X = E(X) =\(\large \sum x P(x)=..?\)

OpenStudy (anonymous):

so say x = 23 what do i do after

OpenStudy (amistre64):

isnt that the same as a weighted average?

OpenStudy (anonymous):

no idea

hartnn (hartnn):

you actually sum the product of X and its probability P(X) for each value of X \(\sum x P(x) = 23*0.16+25*0.09+...+38*0.21=...?\)

OpenStudy (anonymous):

oh now i see! thank you!

hartnn (hartnn):

welcome ^_^

OpenStudy (amistre64):

hmm, i spose since the Xs dont add to 100 that my wghtd avg goes awry :/

hartnn (hartnn):

Xs don't *need* to add upto 100 ... P(X) needs to add upto 1....

OpenStudy (amistre64):

yeah .... now im wondering where i misplaced my wghtd avg :) granted I know the E(X) is the sum like you posted. im going to have to go stab my ear with a pencil now